Turkish vocabulary

David Pierce

November 28, 2009

1 aban- lean against 2 abart- exaggerate 3 acemi inexperienced 4 acı- hurt, pity 5 acın- grieve 6 aciz, -czi inability 7 ada- devote 8 adap rules of good manners 9 adım step 10 ağar- whiten 11 ahır stable, shed 12 âhir last, final 13 ak- flow 14 akciğer lung 15 akın raid 16 akis, -ksi reflexion 17 aksa- limp 18 aktar- turn over 19 aleyh against 20 algı perception 21 âmir commanding 22 an- think of 23 anımsa- remember 24 apar- carry away 25 ar- be exhausted 26 art back 27 art- increase cf. yüksel-28 asgari minimum 29 aş- cross 30 ayır- separate 31 ayrı ayrı individually = tek tek 32 ayrıcalık privilege 33 azal- become less 34 azınlık minority cf. çoğunluk 35 bağır bosom 36 bakım care 37 baskı oppression 38 belirt- make clear 39 bertaraf setting aside 40 beyan, beyanname declaration 41 biçim exposed 42 bilinç consciousness 43 birey individual 44 birim unit 45 bizzat personally 46 bol plentiful 47 boynuz horn 48 boyut dimension 49 bunaltı discomfort, anxiety 50 camia group 51 cesaret bravery 52 cesur brave = yiğit 53 cinayet murder 54 çal- play, steal 55 çarçur squandering 56 çarp- hit, strike; multiply 57 çekiş- quarrel 58 çerçeve frame 59 çevir- turn, surround 60 çim lawn 61 çoğunluk majority cf. azınlık 62 çoluk çocuk household, children 63 dağ mountain 64 dağ brand 65 dağla- brand 66 dahi also, even 67 dâhi genius 68 dava suit, trial 69 davran- bestir 70 dayan- (-a) lean on, be supported by 71 değnek stick, cane 72 dek ≈ kadar 73 denetim control 74 denge balance 75 derle- collect 76 deyiş style of speech 77 dilen- beg 78 diren- resist 79 direniş resistance 80 dit- shred 81 dize verse 82 doku tissue 83 dolaş- wander 84 dolayı because of 85 duman smoke 86 durum situation 87 duyar sensitive 88 duyu sense 89 düdük whistle 90 düşüş fall, decrease 91 dürüst correct 92 ecnebi stranger, foreigner 93 edin- get, acquire 94 eğil- be bent 95 eksen axis 96 eleştir- criticize 97 emin certain (person) cf. kesin 98 ere- melt 99 erte next, following 100 ertele- postpone 101 esas foundation, fundamental 102 eser work, opus 103 etki effect 104 ezgi melody 105 fayda profit, use = yarar 106 felaket disaster 107 gelişigüzel at random 108 gereksin- consider necessary 109 gıdım ? 110 görkem pomp 111 gözyaşı tear 112 günah sin 113 güt- herd, cherish 114 hal, -i state, condition 115 hâlâ still 116 harca- spend 117 harman threshing, harvest 118 harmanla- blend 119 hata mistake 120 hattâ even 121 havale assignment, referral 122 hayıf injustice, pity 123 hemfikir like-minded 124 hoşgürü tolerance 125 husus matter, subject 126 ıslak wet cf. yaş 127 ibare wording 128 idare managing 129 iddia claim 130 ifade expression 131 iflas bankrupt 132 ihale awarding contract, auctioning 133 ihbar denunciation 134 ihraç exporting, expulsion 135 ikamet residence 136 ikametgâh place of residence 137 il- tie, baste 138 ilan announcement 139 ileri sür- assert 140 ilerle- go forward 141 ilinek accident 142 iliş- be attached 143 ilişkin concerning 144 ilke principle 145 iltifat treating with friendliness 146 imece collective labor 147 inanç belief 148 incele- examine 149 intihar suicide 150 ise as for 151 ishal diarrhea 152 istilâ invasion 153 isyan rebellion 154 işaret sign 155 işletici operator of a business 156 işletmeci manager 157 iştah appetite 158 itham accusation 159 itina care 160 itiraf confession 161 itiraz protest 162 kaba rude 163 kabar- swell 164 kaç- escape, flee 165 kalıntı residue 166 kamuoyu public opinion 167 kanıt proof 168 kapla- cover 169 kapsam scope 170 karşıla- meet, cover 171 kasıt intention 172 karaciğer liver 173 karakol police station 174 karmaşa chaos 175 kaşe seal implement or impression 176 kat- add 177 kavga fight 178 kavra- comprehend 179 kavram concept 180 kavuş- be reunited with, reach 181 kaydet- register 182 kaygı worry 183 kayıt registration 184 kazanç gain, earnings 185 kesin certain (thing) cf. emin 186 kırık crack(ed) 187 kıs- reduce 188 kısım part, piece 189 kışkırt- incite 190 kıvrım curl 191 kızgın red-hot, angry, in heat 192 kıyas comparison 193 kibar noble, polite cf. nazik 194 kitle mass 195 komutan commander 196 konum position, location 197 korsan pirate 198 koru- protect 199 kucak embrace, lap 200 kurum institution 201 kus- vomit 202 kuşak generation, belt 203 kuşat- surround 204 kuşkusuz doubtless 205 lâf empty words 206 mahkûm condemned 207 mahsus -a peculiar to 208 maliyet cost 209 maruz exposed 210 mazhar (-a) object (of honors, &c.) 211 mecbur(en) compulsori(ly) 212 mekân place 213 meşgul busy 214 mezbaha slaughterhouse 215 miras inheritance 216 muhtemel(en) probabl(y) 217 munzam ? 218 mutlak absolute 219 mutlaka certainly 220 müdahale interference 221 mülkiyet possession 222 müsaade permission 223 müteahhit contractor 224 müttefik ally 225 müzakere discussion 226 name document, &c. in compounds 227 nazik delicate, polite cf. kibar 228 nice how many 229 nitekim likewise 230 nitele- qualify 231 nitelik quality 232 niyet intention 233 olanak possibility 234 olay event 235 olumsuz negative 236 onay approval 237 ordu army 238 ortam environment 239 oyunbaz playful 240 ozan folk poet cf. şair 241 öbek group, mass 242 öneri suggestion, proposal 243 önle- prevent 244 önlem precaution 245 ör- knit 246 örgüt organization 247 ötürü because of 248 özen care, pains 249 özgün original 250 paha value 251 pansuman wound dressing 252 perhiz fasting (Christian or Jewish) 253 rağmen despite 254 rant unearned income 255 ruhsat authorization 256 sağla- provide 257 sağla- move right 258 saha field 259 sahip çık- be protector 260 sakatla- injure 261 sakın- beware 262 sakınca objection, drawback 263 sakin quiet; dweller 264 salgın epidemic 265 san fame, title 266 san- suppose 267 sanık suspect, accused 268 sanki as if 269 sarmaşık ivy, vine 270 satır line of writing 271 satır meat cleaver 272 sayesinde owing to positively; cf. yüzünden 273 saygın respected 274 sermaye capital 275 sesle- hearken 276 seslen- call out 277 seyyar itinerant 278 sığ- fit into 279 sız- leak 280 sol- fade, wilt 281 solu- breathe, pant 282 solunum respiration 283 sorumlu responsible 284 soy- rob 285 söylenti rumor 286 subay officer neol. 287 sus- be(come) quiet 288 suskun quiet 289 sürece while 290 süreç process 291 sürükle- drift 292 şair poet cf. ozan 293 şantaj blackmail 294 şaş- be amazed 295 şaşır- be at a loss 296 şebeke network, system 297 şemsiye umbrella 298 şer evil 299 şerh explanation 300 şırıl şırıl with a gentle continuous plash 301 şikâyet complaint 302 şöhret reputation 303 şura council 304 tabaka layer, sheet 305 taciz annoyance 306 takdir appreciation 307 takip pursuit 308 talep demand 309 tamamen completely 310 tanım definition 311 tartış- struggle, dispute 312 tasarı proposal, bill 313 taşı- carry 314 tavır attitude 315 tavsiye recommendation 316 tehdit threat 317 tek tek individually = ayrı ayrı 318 teklif proposal 319 tekme kick 320 tekmele- kick 321 temenni heartfelt wish 322 temsil representation 323 teslim delivering, yielding 324 tezgâh counter 325 tercih preference = seçim 326 ters reverse 327 tescil approval = onay 328 tetik quick, trigger 329 tını tone 330 tiran tyrant 331 tutuk hesitant; under arrest 332 tutum manner 333 tüket- consume 334 tür kind, species 335 türe- spring up 336 uğra- stop by, undergo 337 uğraş struggle 338 ulaş- reach, arrive at 339 unsur element = öğe 340 unvan title 341 ut, utanç shame, modesty 342 uyanık awake 343 uyruk citizen 344 uzantı extension, prolongation 345 uzman expert 346 ürk- be frightened 347 üslup style 348 vade term, period of time 349 vana valve, stopcock 350 vazgeç- give up 351 vaziyet condition, situation 352 vekil agent &c. 353 veryansın et- attack verbally 354 vurgu stress, accent 355 yak- light, burn 356 yakala- catch 357 yakın near 358 yakın- (-dan) complain (about) 359 yakış- befit 360 yan- burn 361 yandaş supporter, follower 362 yanıt answer 363 yansı- reflect 364 yarala- wound 365 yarar benefit 366 yarat- create 367 yargı idea, judgment 368 yarış race 369 yaş wet, fresh cf. ıslak, gözyaşı 370 yay- spread 371 yekûn total amount 372 yemin oath 373 yerleş- settle 374 yet- suffice 375 yetişkin mature 376 yetki authority 377 yık- pull down 378 yıpran- wear out 379 yit- be lost = kaybol-380 yitir- lose = kaybet-381 yitik loss = kayıp 382 yöntem method 383 yumak ball of string 384 yut- swallow 385 yüksel- rise cf. art-386 yürek heart = kalp 387 yürürlük validity, being in force 388 yüzünden because of negatively; cf. sayesinde 389 zabıta municipal police 390 zan opinion 391 zannet- suppose = san-392 zarif elegant 393 zaten anyway, in fact 394 zehir poison 395 zeki intelligent 396 zevk taste 397 zihin, -hni mind 398 zihniyet, -ti mentality 399 ziyaret visit 400 zorunlu necessary